(Calculating rates of return) Balxo Balloons manufactures and distributes birthday balloons. At the beginning of the year Blaxo's common stock was selling for $22.99 but by year end it was only $20.14. If the firm paid a total cash dividend of $2.41 during the year, what rate of return would you have earned if you had purchased the stock exactly one year ago? What would your rate of return have been if the firm had paid no cash dividend?
Answer #1

a) Rate of return =(Ending price-beg price+dividend)/beg =(20.14-22.99+2.41)/22.99 = -1.91%

b) If no dividend =(20.14-22.99)/22.99 = -12.40%